Find the area bounded by the curve y = x^3-12 x and the x-axis. Sketch the curve indicating the area bounded.
1
Expert's answer
2020-12-20T16:54:06-0500

Solution. Find the intersection points of the lines


"y=x^3-12x""y=0"

Therefore get "x^3-12x=0"

"x(x^2-12)=0"

The roots of the equation are


"x_1=0""x_2=\sqrt{12}=2\sqrt{3}""x_3=-\sqrt{12}=-2\sqrt{3}"

Sketch the curve indicating the area bounded.


Find the area bounded by the curve y = x^3-12 x and the x-axis. (considering the symmetry of the figure)


"A=2\int_{-2\sqrt{3}}^0(x^3-12x-0)dx=\int_{-2\sqrt{3}}^0(2x^3-24x)dx="

"=(\frac{x^4}{2}-12x^2)|_{-2\sqrt{3}}^0=0-72+144=72"

Answer. 72.